On January 1, a company issued and sold a $450,000, 3%, 10-year bond payable, and received proceeds of $444,000. Interest is payable each June 30 and December 31. The company uses the straight-line method to amortize the discount. The carrying value of the bonds immediately after the first interest payment is:

Multiple Choice

  • $450,000.

  • $449,700.

  • $450,300.

  • $443,700.

  • $444,300.
Answer #1

discount on issue of bond = 450000 - 444000

= $ 6000

Cash interest paid on 30th June = face value of bond* coupon rate

= 450000*3%*1/2

= 13500*1/2

= $ 6750

Carrying value at the time of issue of bond = 444000

Interest to be amortized per half year under straight line method = discount on issue of bond/(no. of years*2)

= 6000/(10*2)

= 6000/20

= $ 300

Interest expenses for the first half year = cash paid interest + bond discount

= 6750 + 300

= 7050

Carrying value of bond immediately after the first interest payment = carrying value of bond at the time of issue + Amortized interest

= 444000 + 300

= $ 444300

so option (d) should be the right answer.

please check with your answer and let me know.